Calculate the incidence rate.

# of new cases per time period: 1067

Population: 8,326,289

Formula: (number of new cases in a time period/population) X K
Answer #1

IR= (number of new cases in a time period/population) X K

IR= (1067/8,326,289) *100,000

= 1.28 * 10^-4 * 10^5 =12.8 ; This means that 12.8 people in every 100,000 will have the disease, or we can also say that 1 person in every 10,000 will have the disease.